# Understanding Code Understandability Improvements in Code Reviews

### v1

@@ 4.3 How likely are understandability improvement comments to be accepted? (RQ3)
We verified whether code review comments that include suggestions for improving code understandability are more likely to be accepted by developers than comments that do not include such suggestions. Our analysis revealed that code review comments that suggest understandability improvements are more likely to be accepted than comments that do not address understandability; the odds ratio is 2.69, with a confidence interval (1.71, 4.22) for a 99% confidence level. We also performed a Chi-Square test to evaluate the relationship between these two variables and found a significant correlation between them. In fact, suggestions for improving code understandability were found to be significantly more likely to be accepted than comments unrelated to understandability ($\chi^{2}(1,N=440)=32.9,p<0.01$).
@@
TABLE IV: What are the odds of the code review suggestion being accepted, given that it is a code understandability improvement?
@@ 4.4 What code changes are found in understandability improvements? (RQ4)
Complex, long, or inadequate logic. 54 patches extracted, moved, or replaced complex, lengthy, or inconsistent code in an attempt to simplify it. Among these, five patches extracted parts of a method to reduce its length (e.g., extracting a method). Also, developers applied 16 patches that replaced an usage of an external API or adopted an external API to simplify complex code, such as replacing a chain of method invocations by an unique API method invocation [key=1700]. In 14 other scenarios, developers replaced expressions in source code by an alternative that better communicates the functionality of the code, e.g., replacing boolean expression by a call to method isEmpty() [key=626]. Additionally, developers applied patches that replace the use of a construct in eight cases, such as replacing an anonymous inner class by a lambda [key=1971] or replacing a lambda with conditional by an if statement [key=1892]. Finally, the developers applied five patches to organize the use of APIs in a source code file, such as expanding imports instead of using wildcards [key=2203], and adding an import statement to avoid the use of the fully-qualified name of a class ( TileSourceManager.TileSourceTemplate template $\rightarrow$ import ...TileSourceManager; private TileSourceTemplate template [key=1779]).
TABLE V: Understandability smell categories and the improvement patches found in the 323 accepted code reviews. The patches with three or more instances are marked with the symbol $\clubsuit$. In the PDF version, each category links to an example.
@@
Inconsistent or disrupted formatting. 37 patches organized and standardized code formatting. These patches manage the use of formatting characters such as space and braces, as well as formatting styles. In 26 of the patches, the developers added space between different code elements (e.g., horizontal space [key=493] and vertical space [key=2050]); and removed space between related code statements (e.g., for declaration and first statement of the block [key=1660]) or extra space between code elements in the same expression (e.g., between the point and the invoked method in an expression [key=1754]). Additionally, developers added other structuring characters (e.g., braces and parentheses) to highlight the region of a code element (e.g., block [key=1078] or operand [key=1453]). Finally, in eight cases developers changed the order or moved code elements aiming to apply a formatting style, such as keeping the code elements in their own lines (e.g., body of if statement in its own line [key=2024]) and arranging method declarations [key=1860].
@@
Answer to RQ4. We found 323 patches where developers added, removed, changed, replaced, and moved code elements to improve code understandability. The most common type of understandability improvement was modifying an identifier to express the meaning of a code element (50). This was followed by removing unused constants, imports, classes, methods, or variables (19), changing text in existing Javadoc to improve grammar or fix a typo (15), and removing duplicate code or processing (13). Implications. These results show that there is considerable variation in understandability improvements; some can be easily supported by linters (e.g., removing unused elements) whereas others require approaches that may be challenging even considering recent advances in machine learning (e.g., removing functionally equivalent code that is not textually similar). They emphasize how diverse tools can be employed to address the different understandability smells.
@@ 4.5 To what extent are accepted code understandability improvements reverted? (RQ5)
We also evaluate whether the (314) patches implementing suggested understandability improvements and integrated into codebase were reverted at some point in the history of the codebase, after the original pull request and the associated code review were closed. Out of the total 314 patches, 82.8% (260/314) are present in the latest version of the codebase, although in 20 cases, the developers renamed or moved the patched files. However, we cannot find the patch in the last version of the codebase in 54 patches. Among these, the developers replaced the patch with an alternative solution in ten observations. For instance, a subsequent patch changed a snippet with functional implementation (commit a0c1cbf9746ed9486f08ce4aa2376828e846c786) by an alternative imperative implementation [key=1453]. In 24 patches, the developers removed the code snippet that contained the patch, e.g., the patch included the deprecated annotation and deprecated methods were deleted [key=2263]. Still, in 18 patches, the developers deleted the entire patched file, e.g., AssemblyJsonServlet.java deleted [key=1935]. Finally, the patches were reverted and kept reverted until the latest version of the codebase in two of these 54 observations. One such case involved the removal of an unnecessary class [key=2073], but the file was added back into the codebase and later renamed.
@@
. Patch status (# analysed patches) Yes No Patch merged (323) 314 (97.2%) 9 (2.8%) Patch reverted in the pull request (9) 0 (0%) 9 (100%) Patch in the last version of the file (314) 260 (82.8%) 54 (17.2%) Patch reverted in the codebase history (54) 2 (3.7%) 52 (96.3%)
@@ 4.6 Do linters contain rules to detect the identified code understandability smells? (RQ6)
Our analysis indicates that 89 out of 323 occurrences of understandability smells can be detected by any of the linters. This highlights wasted human effort, as automated tools could be pointing these issues out. On the one hand, Checkstyle and Sonarqube are responsible for detecting more than half of detectable occurrences. On the other hand, Spotbugs detects only seven occurrences overall, where six are related to unnecessary code. Furthermore, we did not find rules that could to detect the occurrences related to Wrong, missing, or inadequate string expression or literal, either because the tools do not inspect string literals or because they are natural language text.
@@
The two most popular understandability smells, Bad identifier and Incomplete or inadequate code documentation, are the second and fourth less detectable by linters, respectively. The majority of the occurrences of these understandability smells are associated to natural language text (e.g., typo in identifier and issue in content of Javadoc). The cases of Bad identifier detected by the linters are those related to style (e.g., parameters starting with capital letter [key=579]). In these cases, rules such as FormalParameterNamingConventions (PMD), Nm: Field names should start with a lower case letter (NM_FIELD_NAMING_CONVENTION) (Spotbugs), and LocalVariableName (Checkstyle) can be used to detect them. In the case of Incomplete or inadequate code documentation, there are only rules for missing documentation, e.g., CommentRequired (PMD); missing or incorrect order of tags in Javadoc, e.g., WriteTag (Checkstyle); and TODO code comments, e.g., Track uses of “TODO” tags (SonarQube). Many of the instances of these smells require in-depth understanding of the associated elements, e.g., to produce an identifier that better represents the purpose of a method or to improve an explanation about an element. At the same time, the most common change applied to Javadoc documentation in our study is fixing typos (Table V), which could be automated.
@@
Among the remaining 234 out of 323 instances of understandability smells, we identify occurrences that could be covered by the investigated linters by performing small adaptations to existing rules. For example, the linters could cover the missing parentheses (e.g., key=1453) by creating an inverse rule to the UselessParentheses rule of PMD or “UnnecessaryParentheses” rule of Checkstyle. This new rule could suggest the inclusion of additional parentheses in boolean expressions to highlight evaluation order. Additionally, the linters could identify opportunities to use primitive types instead of the corresponding boxed types (e.g., key=1201). As another example, the linters could detect when a exception is created without a message (e.g., key=1345). However, other instances of understandability smells would require more sophisticated analyses. For example, to detect the occurrences of incomplete documentation, unnecessary logging messages, confusing string messages, inadequate use of method call chains, and missing logging, it would be necessary to understand the context of the source code, the culture of the project, the intention of the developers, predict the system behavior at run time, or to work with both Java and natural languages.
@@
Answer to RQ6. We found 89 out of 323 occurrences of understandability smells that could be detect by the linters. Some of the remaining 234 occurrences of understandability smells could be detected by creating new rules in these linters, while other occurrences require a deeper and context-dependent analysis that goes beyond what linters typically do. Implications. The four linters combined have rules to cover less than 30% of the instances of understandability smells from our dataset. This suggests that developers waste effort during code reviews by alluding to issues that existing automated tools could detect. Among the cases that the linters cannot detect, many of them require an in-depth “understanding” of the program under analysis or the context in which it is inserted. In particular, current linters are unable to cope with code understandability issues related to natural language. Some of these cases present opportunities to devise new tools, e.g., to better summarize the intent of a code element or to give it a more illustrative name. Others could be addressed by simple improvements to current linters, e.g., pointing out typos in documentation, identifiers, and string literals. This also applies to rules related to other aspects, such as messages in exceptions and the use (or not) of boxed types.
@@ 5 Threats to validity
Selection of projects. We constructed our dataset by mining software repositories from GitHub.However, researchers must be cautious about potential pitfalls [37]. To ensure data quality, we applied filters to exclude personal projects, repositories with small contributor groups, and forked repositories. Moreover, we utilized a heuristic to select projects with an active code review process during a specific period. Also, we manually analyzed each remaining repository to guarantee the relevance of the dataset, excluding non-software repositories. Additionally, it is possible that some projects may have configured an ASAT (Automated Static Analysis Tool) to prevent certain issues related to code understandability from being introduced into the codebase. However, since we did not assess whether any project adopted this approach, it is possible that some understandability smells may not have been identified by our study
Size of the sample of code review comments. Before analyzing our sample of 2,401 code review comments to determine whether they are about code understandability improvements, we explored various natural language processing techniques, such as LDA, GSDMM, LSA, and LDA Guided, to extract topics and create clusters for both understandability and non-understandability improvements. Unfortunately, we encountered challenges in obtaining satisfactory results due to two main reasons: (i) our training sample was relatively small, both in terms of the number of comment instances and the length of the text; and (ii) there were significant similarities between comments related and unrelated to code understandability. This problem has been reported by previous work as well [68, 69].
@@
We also conducted a manual analysis of potential rules in four linters that could detect issues highlighted by reviewers. For each instance of an understandability smell that we analyzed, we searched through a set of 1,315 rules to find matching ones that could identify the issue. Some of these rules can be configured to detect the problem, and we considered all possible configurations. However, we decided not to include rules that relied on regex patterns, due to their specificity. It’s important to note that there may be some rules that we missed during this process, which could result in some issues being classified as not covered by linters.
@@ 6 Related Work
In this section we examine related work. We start by contrasting the understandability smells and improvements we have identified in practice with recommendations and experimental results reported in the literature (Section 6.1). This helps us identify mismatches and accordance between research and practice, as well as research gaps. We then review previous work discussing models to assess code understandability (Section 6.2), following up with papers that have investigated manifestations of confusion and reactions to lack of necessary information during code reviews (Section 6.3).
@@
Our focus in this paper is on understandability improvements, i.e., situations where existing code exhibits an understandability smell and is replaced by an alternative solution that is more understandable. Thus, in this section we focus on the five understandability smells where, to the best of our knowledge, alternative solutions were compared in existing studies from a human-centric perspective, e.g., based on the preference or performance of study participants: Bad identifier, Unnecessary code, Inconsistent or disrupted formatting, Complex, long, or inadequate logic, and Missing constant usage.
Bad identifier. Chaudhary and Sahasrabuddhe [71] evaluated the impact on code understandability of using identifier names unrelated to the program domain. This study concluded that the identifier names with meaning related to the problem domain facilitate understanding. In this direction, the practitioners improve the identifier names to better match their intent and type. Santos and Gerosa [5] compared using fully-qualified names and import clauses with shorter type names. They found out that participants preferred to avoid fully-qualified names and add import clauses, which is what the practitioners in our study prefer. Schankin et al. [72] investigated whether a more descriptive name (e.g., singleWordParts) is easier to read than a single-word one (e.g., parts). They found that a descriptive name was the best in one scenario, and there was no difference in the other. The practitioners also considered improving understandability with descriptive names. Sharif and Maletic [73] and Binkley et al. [74] investigated the effect of the style of the identifier (underscore and camel case) on code understandability. Sharif and Maletic found that the use of underscore to separate parts of an identifier improves understandability, whereas Binkley et al. found out that camel case works better. For practitioners in Java, using the camel case style for non-constant identifiers improves code understandability.
Unnecessary code. Gopstein et al. [18] and Torres et al. [35] investigated whether the presence of dead, unreachable, or repeated code (e.g., v1 = 1; v1 = 2), leads to misunderstanding in C and JavaScript programs, respectively. Both studies found out that the presence or absence of this pattern did not significantly impact code understandability. In contrast, practitioners seem to believe this is an important issue in practice: 13.3% of all the 323 analyzed code review comments pertain to unnecessary code. The practitioners removed the unnecessary code, commented out source code, and duplicated code to improve code understandability.
@@
Complex, long, or inadequate logic. Gopstein et al. [18], Medeiros et al. [7], Torres et al. [35], Costa et al. [78], and Langhout and Aniche [9] investigated which ones promotes better understandability: the conditional operator or if statements, in C, JavaScript, Python, and Java small programs. Gopstein et al. and Medeiros et al. found that if statements improve code understandability (in C programs) whereas Langhout and Aniche, Costa et al., and Torres et al. did not find a statistically significant difference. The practitioners considered both if statement and ternary operator to improve code understandability in distinct scenarios. Lucas et al. [79] compared the use of lambdas and anonymous inner classes based on the opinions of 28 experienced Java developers. Their results are aligned with our findings when examining code review comments: practitioners prefer lambda expressions. Most of the instances of the Complex, long, or inadequate understandability smell are either too specific to be studied in a more general context, e.g., object creation that can be avoided by using an existing attribute, or, to the best of our knowledge, have not been studied, e.g., using try statements that manage resources, instead of try with resources.

### v2

@@ 4.3 How likely are understandability improvement comments to be accepted? (RQ3)
We verified whether code review comments that include suggestions for improving code understandability are more likely to be accepted by developers than comments that do not include such suggestions. Our analysis revealed that code review comments that suggest understandability improvements are more likely to be accepted than comments that do not address understandability: the odds ratio is 2.69, with a confidence interval [1.71, 4.22] for a 99% confidence level. We also performed a Chi-Square test to evaluate the relationship between these two variables and found a significant correlation between them. In fact, suggestions for improving code understandability were found to be significantly more likely to be accepted than comments unrelated to understandability ($\chi^{2}(1,N=440)=32.9,p<0.01$).
@@
TABLE IV: Frequency of accepted vs. not accepted code review suggestions with and without understandability improvements
@@ 4.4 What code changes are found in understandability improvements? (RQ4)
Complex, long, or inadequate logic. 54 patches extracted, moved, or replaced complex, lengthy, or inconsistent code in an attempt to simplify it. Among these, five patches extracted parts of a method to reduce its length (e.g., extracting a method). Also, developers applied 16 patches that replaced a usage of an external API or adopted an external API to simplify complex code, such as replacing a chain of method invocations by a unique API method invocation [key=1700]. In 14 other scenarios, developers replaced expressions in source code by an alternative that better communicates the functionality of the code, e.g., replacing boolean expression by a call to method isEmpty() [key=626]. Additionally, developers applied patches that replace the use of a construct in eight cases, such as replacing an anonymous inner class by a lambda [key=1971] or replacing a lambda with conditional by an if statement [key=1892]. Finally, the developers applied five patches to organize the use of APIs in a source code file, such as expanding imports instead of using wildcards [key=2203], and adding an import statement to avoid the use of the fully-qualified name of a class ( TileSourceManager.TileSourceTemplate template $\rightarrow$ import ...TileSourceManager; private TileSourceTemplate template [key=1779]).
TABLE V: Understandability smell categories and the improvement patches found in the 323 accepted code reviews. The patches with three or more instances are marked with the symbol $\clubsuit$. In the PDF version, each patch links to an example.
@@
Inconsistent or disrupted formatting. 37 patches organized and standardized code formatting. These patches manage the use of formatting characters such as space and braces, as well as formatting styles. In 26 of the patches, the developers added space between different code elements (e.g., horizontal space [key=493] and vertical space [key=2050]); and removed space between related code statements (e.g., for declaration and the first statement of the block [key=1660]) or extra space between code elements in the same expression (e.g., between the point and the invoked method in an expression [key=1754]). Additionally, developers added other structuring characters (e.g., braces and parentheses) to highlight the region of a code element (e.g., block [key=1078] or operand [key=1453]). Finally, in eight cases developers changed the order or moved code elements aiming to apply a formatting style, such as keeping the code elements in their own lines (e.g., body of if statement in its own line [key=2024]) and arranging method declarations [key=1860].
@@
Answer to RQ4. We found 323 patches where developers added, removed, changed, replaced, and moved code elements to improve code understandability. The most common type of understandability improvement was modifying an identifier to express the meaning of a code element (50). This was followed by removing unused constants, imports, classes, methods, or variables (19), changing text in existing Javadoc to improve grammar or fix a typo (15), and removing duplicate code or processing (13). Implications. These results show that there is considerable variation in understandability improvements: some can be easily supported by linters (e.g., removing unused elements) whereas others require approaches that may be challenging even considering recent advances in machine learning (e.g., removing functionally equivalent code that is not textually similar). They emphasize how diverse tools can be employed to address the different understandability smells.
@@ 4.5 To what extent are accepted code understandability improvements reverted? (RQ5)
We also evaluate whether the (314) patches implementing suggested understandability improvements and integrated into the codebase were reverted at some point in the history of the project after the original pull request and the associated code review were closed. Out of the total 314 patches, 82.8% (260/314) are present in the latest version of the codebase, although in 20 cases, the developers renamed or moved the patched files. However, we cannot find the patch in the last version of the codebase in 54 patches. Among these, the developers replaced the patch with an alternative solution in ten observations. For instance, a subsequent patch changed a snippet with functional implementation (commit a0c1cbf9746ed9486f08ce4aa2376828e846c786) by an alternative imperative implementation [key=1453]. In 24 patches, the developers removed the code snippet that contained the patch, e.g., the patch included the deprecated annotation and deprecated methods were deleted [key=2263]. Still, in 18 patches, the developers deleted the entire patched file, e.g., AssemblyJsonServlet.java deleted [key=1935]. Finally, the patches were reverted and kept reverted until the latest version of the codebase in two of these 54 observations. One such case involved the removal of an unnecessary class [key=2073], but the file was added back into the codebase and later renamed.
@@
. Patch status (# analyzed patches) Yes No Patch merged (323) 314 (97.2%) 9   (2.8%) Patch reverted in the pull request (9) 0     (0%) 9   (100%) Patch in the last version of the file (314) 260 (82.8%) 54 (17.2%) Patch reverted in the codebase history (54) 2   (3.7%) 52 (96.3%)
@@ 4.6 Do linters contain rules to detect the identified code understandability smells? (RQ6)
Our analysis indicates that 89 out of 323 occurrences of understandability smells can be detected by any of the linters. This highlights wasted human effort, as automated tools could be pointing these issues out. On the one hand, Checkstyle and Sonarqube are responsible for detecting more than half of detectable occurrences. On the other hand, Spotbugs detects only seven occurrences overall, where six are related to unnecessary code. Furthermore, we did not find rules that could detect the occurrences related to Wrong, missing, or inadequate string expression or literal, either because the tools do not inspect string literals or because they are natural language text.
@@
The two most popular understandability smells, Bad identifier and Incomplete or inadequate code documentation, are the second and fourth less detectable by linters, respectively. The majority of the occurrences of these understandability smells are associated with natural language text (e.g., typo in identifier and issue in content of Javadoc). The cases of Bad identifier detected by the linters are those related to style (e.g., parameters starting with capital letter [key=579]). In these cases, rules such as FormalParameterNamingConventions (PMD), Nm: Field names should start with a lower case letter (NM_FIELD_NAMING_CONVENTION) (Spotbugs), and LocalVariableName (Checkstyle) can be used to detect them. In the case of Incomplete or inadequate code documentation, there are only rules for missing documentation, e.g., CommentRequired (PMD); missing or incorrect order of tags in Javadoc, e.g., WriteTag (Checkstyle); and TODO code comments, e.g., Track uses of “TODO” tags (SonarQube). Many instances of these smells require an in-depth understanding of the associated elements, e.g., to produce an identifier that better represents the purpose of a method or to improve an explanation about an element. At the same time, the most common change applied to Javadoc documentation in our study is fixing typos (Table V), which could be automated.
@@
Among the remaining 234 out of 323 instances of understandability smells, we identify occurrences that could be covered by the investigated linters by performing small adaptations to existing rules. For example, the linters could cover the missing parentheses (e.g., key=1453) by creating an inverse rule to the UselessParentheses rule of PMD or “UnnecessaryParentheses” rule of Checkstyle. This new rule could suggest the inclusion of additional parentheses in boolean expressions to highlight evaluation order. Additionally, the linters could identify opportunities to use primitive types instead of the corresponding boxed types (e.g., key=1201). As another example, the linters could detect when an exception is created without a message (e.g., key=1345). However, other instances of understandability smells would require more sophisticated analyses. For example, to detect the occurrences of incomplete documentation, unnecessary logging messages, confusing string messages, inadequate use of method call chains, and missing logging, it would be necessary to understand the context of the source code, the culture of the project, the intention of the developers, predict the system behavior at run time, or to work with both Java and natural languages.
@@
Answer to RQ6. We found 89 out of 323 occurrences of understandability smells that could be detected by the linters. Some of the remaining 234 occurrences of understandability smells could be detected by creating new rules in these linters, while other occurrences require a deeper and context-dependent analysis that goes beyond what linters typically do. Implications. The four linters combined have rules to cover less than 30% of the instances of understandability smells from our dataset. This suggests that developers waste effort during code reviews by alluding to issues that existing automated tools could detect. Among the cases that the linters cannot detect, many of them require an in-depth “understanding” of the program under analysis or the context in which it is inserted. In particular, current linters are unable to cope with code understandability issues related to natural language. Some of these cases present opportunities to devise new tools, e.g., to better summarize the intent of a code element or to give it a more illustrative name. Others could be addressed by simple improvements to current linters, e.g., pointing out typos in documentation, identifiers, and string literals. This also applies to rules related to other aspects, such as messages in exceptions and the use (or not) of boxed types.
@@ 5 Threats to validity
Selection of projects. We constructed our dataset by mining software repositories from GitHub. However, researchers must be cautious about potential pitfalls [37]. To ensure data quality, we applied filters to exclude personal projects, repositories with small contributor groups, and forked repositories. Moreover, we utilized a heuristic to select projects with an active code review process during a specific period. Also, we manually analyzed each remaining repository to guarantee the relevance of the dataset, excluding non-software repositories. Additionally, some projects may have configured an ASAT (Automated Static Analysis Tool) to prevent certain issues related to code understandability from being introduced into the codebase. However, since we did not assess whether any project adopted this approach, it is possible that some understandability smells may not have been identified by our study.
Size of the sample of code review comments. Before analyzing our sample of 2,401 code review comments to determine whether they are about code understandability improvements, we explored various natural language processing techniques, such as LDA, GSDMM, LSA, and LDA Guided, to extract topics and create clusters for both understandability and non-understandability improvements. Unfortunately, we encountered challenges in obtaining satisfactory results due to two main reasons: (i) our training sample was relatively small, both in terms of the number of comment instances and the length of their text, and (ii) there were significant similarities between comments related and unrelated to code understandability. This problem has been reported by previous work as well [68, 69].
@@
We also conducted a manual analysis of rules in four linters that could potentially detect understandability smells highlighted by reviewers. For each instance of an understandability smell that we analyzed, we searched through a set of 1,315 rules to find the ones that could identify the smell. Some of these rules can be configured to detect a smell, and we considered all possible configurations. However, we decided not to include rules that relied on regex patterns due to their specificity. It is important to note that there may be some rules that we missed during this process, which could result in some issues being classified as not covered by linters.
@@ 6 Related Work
In this section, we examine related work. We start by contrasting the understandability smells and improvements we have identified in practice with recommendations and experimental results reported in the literature (Section 6.1). This helps us identify mismatches and accordance between research and practice, as well as research gaps. We then review previous work discussing models to assess code understandability (Section 6.2), following up with papers that have investigated manifestations of confusion and reactions to lack of necessary information during code reviews (Section 6.3).
@@
Our focus in this paper is on understandability improvements, i.e., situations where existing code exhibits an understandability smell and is replaced by an alternative solution that is more understandable. Thus, in this section, we focus on the five understandability smells where, to the best of our knowledge, alternative solutions were compared in existing studies from a human-centric perspective, e.g., based on the preference or performance of study participants: Bad identifier, Unnecessary code, Inconsistent or disrupted formatting, Complex, long, or inadequate logic, and Missing constant usage.
Bad identifier. Chaudhary and Sahasrabuddhe [71] evaluated the impact on code understandability of using identifier names unrelated to the program domain. This study concluded that the identifier names with meaning related to the problem domain facilitate understanding. In this direction, the practitioners improve the identifier names to better match their intent and type. Santos and Gerosa [5] compared using fully-qualified names and import clauses with shorter type names. They found out that participants preferred to avoid fully-qualified names and add import clauses, which is what the practitioners in our study prefer. Schankin et al. [72] investigated whether a more descriptive name (e.g., singleWordParts) is easier to read than a single-word one (e.g., parts). They found that a descriptive name was the best in one scenario, and there was no difference in the other. The practitioners also considered improving understandability with descriptive names. Sharif and Maletic [73] and Binkley et al. [74] investigated the effect of the style of the identifier (underscore and camel case) on code understandability. Sharif and Maletic found that the use of an underscore to separate parts of an identifier improves understandability, whereas Binkley et al. found out that camel case works better. For practitioners in Java, using the camel case style for non-constant identifiers improves code understandability.
Unnecessary code. Gopstein et al. [18] and Torres et al. [35] investigated whether the presence of dead, unreachable, or repeated code (e.g., v1 = 1; v1 = 2), leads to misunderstanding in C and JavaScript programs, respectively. Both studies found that the presence or absence of this pattern did not significantly impact code understandability. In contrast, practitioners seem to believe this is an important issue in practice: 13.3% of all the 323 analyzed code review comments pertain to unnecessary code. The practitioners removed the unnecessary code, commented out source code, and duplicated code to improve code understandability.
@@
Complex, long, or inadequate logic. Gopstein et al. [18], Medeiros et al. [7], Torres et al. [35], Costa et al. [78], and Langhout and Aniche [9] investigated which one promotes better understandability: the conditional operator or if statements, in C, JavaScript, Python, and Java small programs. Gopstein et al. and Medeiros et al. found that if statements improve code understandability (in C programs) whereas Langhout and Aniche, Costa et al., and Torres et al. did not find a statistically significant difference. The practitioners considered both if statement and ternary operator to improve code understandability in distinct scenarios. Lucas et al. [79] compared the use of lambdas and anonymous inner classes based on the opinions of 28 experienced Java developers. Their results are aligned with our findings when examining code review comments: practitioners prefer lambda expressions. Most of the instances of the Complex, long, or inadequate understandability smell are either too specific to be studied in a more general context, e.g., object creation that can be avoided by using an existing attribute, or, to the best of our knowledge, have not been studied, e.g., using try statements that manage resources, instead of try with resources.